OpenStudy (anonymous):

what quantity of energy does it take to convert .500kg ice at -20c to steam at 250c? specific heat capacities: ice, 2.03 J/gc liquide 4.2Jlgc steam 2.0J/gc, Hvap=40.7kj.mol; Hfus-6.02 kj/mol? Please help me i have a test tom

OpenStudy (nincompoop):

can you try to solve this? so I know where your troubles are

OpenStudy (anonymous):

yeah i can try can you wait for me

OpenStudy (nincompoop):

ye

OpenStudy (anonymous):

okay so first i coverted 500kg to 500g then i multiplied (2.03)(500g)(120)=121800 next i did 500(1/18.02)*6.02=167.03kkj Now im stuck

OpenStudy (nincompoop):

laughing out loud

OpenStudy (nincompoop):

500 kg to 500 g? no

OpenStudy (anonymous):

i dont then

OpenStudy (nincompoop):

okay work on your conversion rates then watch this video http://www.kentchemistry.com/links/Matter/HeatingCurve.htm

OpenStudy (anonymous):

alrighty buddy first of all, 500g doesn't equal 500kg. the conversion rate is that there are 1000 grams in a kilogram so 500kg = 500*1000 g

OpenStudy (anonymous):

oh wow huh okay thank you so much for catching that big mistake

OpenStudy (anonymous):

im using q=s*m*t right?

OpenStudy (anonymous):

wait its not 500 its .500 so the conversion is correct

OpenStudy (anonymous):

i cant get it please help me